OpenStudy (sammietaygreen):

given f(x)=x^2-2x, find f(4)

Nnesha (nnesha):

f(4) means x=4 so replace x with 4

rishavraj (rishavraj):

plug x = 4 in the equation

OpenStudy (anonymous):

f(4) ----> replace in f(x) the x with 4, so it is f(4)=4^2-2*4=16-8=8

Nnesha (nnesha):

you went 2 steps ahead.

OpenStudy (anonymous):

she said '' find f(4) '' not how can i find f(4)

Nnesha (nnesha):

*copy paste* that's exact question from her notes but that doesn't mean we have to solve for them.